OpenStudy (anonymous):

multiply (5-3i)(6+4i) HELP :( im studying for my test and i cant figure it out! my test if friday!

OpenStudy (anonymous):

(5-3i)(6+4i)=5(6+4i)-3i(6+4i)=30+20i-18i-12(i^2)=30+2i-12(-1)=42+2i